In each of the pairwise comparisons for the Bonferroni Multiple Comparison method, what are the degrees of freedom for the t-statistic?A researcher investigated four different word lists for use in hearing assessments. She wanted to know whether the lists were equally difficult to understand in the presence of a noisy background. To find out, she tested 96 subjects with normal hearing. She randomly assigned 24 to each of the four word lists and measured the number of words perceived correctly in the presence of background noise. Here are the summary statistics for number of words perceived correctly for each of the four words lists.A researcher investigated four different word lists for use in hearing assessments. She wanted to know whether the lists were
Answer #1

here total sample size N =24+24+24+24 =96

and number of groups =K =4

therefore degree of freedom for the t statistic =N-k =96-4 =92